OpenStudy (anonymous):

select all that are equivalent to 36x -30 a.18x-15 b.6(6x+5) c.6 x 6x- 6 x 5 d.6x

OpenStudy (cwrw238):

what do you get if you distribute 6 over 6x + 5 ? (option B)

OpenStudy (cwrw238):

can you simplify option c? Do the multiplications first

OpenStudy (anonymous):

i dont think u can simplfy c

OpenStudy (cwrw238):

the way its written is a bit confusing Seems to me that the first and third 'x' s mean multiply.

OpenStudy (anonymous):

yeah its fro multiplying

OpenStudy (cwrw238):

the other is mean the variable x

OpenStudy (anonymous):

yes

OpenStudy (cwrw238):

other x means variable x

OpenStudy (cwrw238):

so what is 6 times 6 x?

OpenStudy (anonymous):

is the x a variable or a multiplication sign

OpenStudy (cwrw238):

thats the variable x

OpenStudy (anonymous):

36x

OpenStudy (cwrw238):

right and 6 times -5?

OpenStudy (anonymous):

times as in the multiplication sign

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

-30

OpenStudy (cwrw238):

ok so now you know...

OpenStudy (anonymous):

what option was that

OpenStudy (anonymous):

so c isnt equivalent

OpenStudy (cwrw238):

put together 36x and -30 and what do you get?

OpenStudy (cwrw238):

6 * 6x - 6 * 5 = 36x - 30

OpenStudy (cwrw238):

write multiply (times) as * then there is no confusion

OpenStudy (anonymous):

6x

OpenStudy (anonymous):

so c and d

OpenStudy (cwrw238):

no its not 6x its c only

OpenStudy (anonymous):

oh okay

OpenStudy (cwrw238):

36x - 30 is not 6x

OpenStudy (anonymous):

then what is it

OpenStudy (cwrw238):

you cannot subtract 30 from 36x because they are unlike terms 36x - 30 is just 36x - 30 It cannot be simplified any further

OpenStudy (anonymous):

ohhhhhh

OpenStudy (cwrw238):

36x - 30x = 6x because they are both like terms

OpenStudy (anonymous):

i get it now thx

OpenStudy (cwrw238):

good yw

OpenStudy (cwrw238):

like 2y - 1 is not y its just 2y - 1 but 2y - y = y

OpenStudy (cwrw238):

- you can only add or subtract LIKE TERMS